OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Human + horse = 74 heads 2(H) + 4(h) = 196 legs

OpenStudy (anonymous):

H = 74 - h 2(74-h) + 4h = 196

OpenStudy (anonymous):

so 148 - 2h + 4h = 196

OpenStudy (anonymous):

2h = 48, h = ....

OpenStudy (anonymous):

24 horses and 50 humans.

OpenStudy (anonymous):

thats what i got :D